1. When 22.0 mL of a 4.07×10-4M sodium sulfide solution is combined with 15.0 mL of a 4.60×10-4M nickel(II) bromide solution does a precipitate form?  (yes or no)

For these conditions the Reaction Quotient, Q, is equal to ____

2. When 15.0 mL of a 6.77×10-4M sodium carbonate solution is combined with 22.0 mL of a 6.39×10-4M lead nitrate solution does a precipitate form?  (yes or no)

For these conditions the Reaction Quotient, Q, is equal to ___
Answer #1

1)

Let us consider a reaction,

Ni+2 + S-2 -----> NiS

Ksp = 13.27 * 10^-11

Q = [Ni+2] [S-2]

[S-2] = M V / Total volume

= 22 *(4.07 * 10 ^-4 )/ (22+15)

= 0.000242

[Ni+2] = M V /Total volume

= (15 * 4.60*10^-4)/ 37

  = 0.000186

Q = [Ni+2][S-2]

= 0.000186 * 0.000242

Q = 4.5 * 10-8

Q << Ksp. Therefore, precipitation does not occurs ( NO)